# Agentic SLMs: Hunting Down Test Smells

### v1

@@ Abstract
Test smells can compromise the reliability of test suites and hinder software maintenance. Although several strategies exist for detecting test smells, few address their removal. Traditional methods often rely on static analysis or machine learning, requiring significant effort and expertise. This study evaluates Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B—small, open language models—for automating the detection and refactoring of test smells through agent-based workflows. We explore workflows with one, two, and four agents across 150 instances of 5 common test smell types extracted from real-world Java projects. Unlike prior approaches, ours is easily extensible to new smells via natural language definitions and generalizes to Python and Golang. All models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%). Analyses were computationally inexpensive and ran efficiently on a consumer-grade hardware. Notably, Phi 4 14B with four agents performed within 5% of proprietary models such as o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. Multi-agent setups outperformed single-agent ones in three out of five test smell types, highlighting their potential to improve software quality with minimal developer effort. For the Assertion Roulette smell, however, a single agent performed better. To assess practical relevance, we submitted 10 pull requests with Phi 4 14B-generated code to open-source projects. Five were merged, one was rejected, and four remain under review, demonstrating the approach’s real-world applicability.
@@ 1 Introduction
Test smells are design flaws in test code that compromise reliability and hinder maintenance [1, 2]. Similar to code smells in production code, they can lead to flakiness, false positives, and false negatives, reducing the effectiveness of test suites. These issues are prevalent in both open-source and industrial settings [3]. Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Numbers[3]. Such smells harm readability and make test evolution more error-prone.
Existing tools for detecting test smells typically rely on static analysis [4, 5] or machine learning [6, 5], but these approaches are often hard to adapt to new smells or other languages [5]. Recent advances in foundation models have opened new possibilities in software engineering [7, 8], yet their use in test smell refactoring—especially through collaborative agentic workflows [9] – remains underexplored.
In this work, we evaluate small, open foundation models – Llama 3.2 3B, Gemma 2 9B, DeepSeek-R1 14B, and Phi 4 14B – for automatically detecting and refactoring test smells through agentic workflows. Our method supports one, two, or four agents and is easily extensible: new smells and refactorings can be defined in natural language. We evaluate 150 instances of 5 frequent test smells from real-world Java projects (Section 3).
Results show that all models detected nearly all test smell instances (pass@5 of 96% with four agents), with Phi 4 14B achieving the highest refactoring accuracy (pass@5 of 75.3%). This performance is within 5% of proprietary models like o1-mini, o3-mini-high, and Gemini 2.5 Pro Experimental using a single agent. Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, a single-agent configuration proved more effective – suggesting that workflow design should consider the specific smell being addressed. To assess real-world impact, we submitted 10 pull requests with Phi 4 14B – generated refactorings to open-source projects. Five were merged, one was rejected, and four remain open, indicating early practical viability.
Preliminary results demonstrate promising generalization across programming languages – specifically Python and Golang – using a the same setup. Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with foundation models to better understand the rationale behind suggested changes. All data and code are publicly available online [10].
@@ 2 Test Smells
Test smells are recurring patterns that reduce test clarity and maintainability. Assertion Roulette [1] occurs when multiple assertions lack descriptive messages, making it hard to trace failures. It can be mitigated by adding messages or splitting checks into separate methods. Conditional Test Logic [11] arises when tests use branching or loops, which may result in unexecuted paths. This can be addressed by isolating each condition in a dedicated test method to ensure full coverage.
Duplicate Assert [3] appears when similar assertions are repeated. Developers should either split assertions into distinct tests if they cover different behaviors or consolidate them to reduce redundancy. Exception Handling is considered a smell when tests use manual try/catch instead of framework features like assertThrows. Replacing manual handling with standard mechanisms improves test readability and intent. Magic Number occurs when numeric literals are hardcoded without explanation, impairing readability. Refactoring involves replacing literals with named constants for better clarity.
@@ 3 Methodology
The primary goal of this study is to evaluate the effectiveness of an agentic approach using small, open models to detect and remove test smells. We analyzed test methods from 11 real-world open-source GitHub projects using JUnit 5 previously studied by Soares et al. [12], including janusgraph, quarkus, testcontainers-java, opengrok, jenkins, lettuce, Mindustry, data-transfer-project, Activiti, flowable-engine, and skywalking. We found that 89% of test cases in these projects contain at most 30 LOC. We limited our analysis to tests of this size. We focused on 5 common test smell types, selecting 30 real-world examples for each. We evaluate four small, open language models – Llama 3.2 3B [13], Gemma 2 9B [14], DeepSeek-R1 14B [15], and Phi 4 14B [16] – using their default configurations. All models were accessed via the Ollama platform and executed locally on a MacBook Pro M3 with 18GB of RAM (January 2025).
Agent communication is managed using the LangChain API [17], and we apply prompting techniques such as Role (Persona) and Chain-of-Thought [18] to enhance reasoning and contextual understanding. The four-agent configuration (Figure 1) distributes responsibilities as follows: Agent 1 detects potential smells; Agent 2 confirms the detection; Agent 3 performs the refactoring; and Agent 4 evaluates the result for correctness and behavior preservation. If there is disagreement, agents engage in an Evaluator-Optimizer loop [9], repeated up to three times.
@@
Agent 1 utilizes the following prompt to detect test smells.
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant with many years of experience that detects test smells.
test_smell
@@
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding expert reviewing the detection of a test smell. Consider the following test smell:
test_smell
@@
It gave the following answer:
explanation
@@
To remove a test smell, Agent 3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent 4 (agent_feedback). {mdframed}[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
test_smell
@@
To evaluate the code proposed by Agent 3, Agent 4 utilizes the following prompt.
{mdframed}
[backgroundcolor=cyan!5, linecolor=black, linewidth=0.5pt] You are a code reviewer specializing in JUnit 5 test smells.
test_smell
@@
The two-agent setup follows the same workflow but omits Agents 2 and 4 – Agent 1 handles detection, and Agent 3 performs the refactoring, without validation or feedback loops. In the single-agent setup, Agent 1 analyzes the test code, identifies potential smells, and applies the corresponding refactoring.
@@ 4 Evaluation
Table I summarizes the detailed performance of the single-agent and multi-agent approaches for detecting and refactoring test smells.
Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama 3.2 3B (Lla), Gemma 2 9B (Gem), DeepSeek-R1 14B (DS) and Phi 4 14B across five test smell types.
@@
Cond. Test Logic 30 3.3% 6.7% 6.7% 0.0% 96.7% 93.3% 100.0% 100.0% 13.3% 0.0% 13.3% 0.0% 96.7% 93.3% 100.0% 100.0% 10% 3.3% 16.7% 3.3%
@@
Exception Handling 30 43.3% 13.3% 40% 36.7% 90.0% 50.0% 93.3% 100.0% 43.3% 23.3% 43.3% 30.0% 86.7% 50.0% 90.0% 100.0% 40.0% 23.3% 40.0% 36.7%
Magic Number 30 73.3% 83.3% 30% 33.3% 93.3% 100.0% 100.0% 100.0% 80.0% 90.0% 43.3% 40.0% 90.0% 100.0% 100.0% 100.0% 70.0% 90.0% 50.0% 33.3%
@@ 4.1 Detecting Test Smells
Gemma 2 9B and Phi 4 14B exhibit good detection performance across most test smells, correctly identifying 98% and 96% instances, respectively, when using two agents. Their performance remains consistent with one or four agents. However, both models fail to detect all instances of the Exception Handling test smell. For the remaining test smells, at least one of these models achieves a perfect detection score. DeepSeek-R1 14B demonstrates the weakest performance, detecting only 78% instances. In contrast, Llama 3.2 3B consistently detects all test smells across all agent configurations (one, two, or four).
We analyze not only the binary yes or no responses but also carefully review the outputs of Agents 1 and 2 to evaluate whether their explanations make sense in detecting test smells. Next, we present examples of outputs generated by Agent 1 when using four agents. For example, Agent 1 correctly identifies a Assertion Roulette test smell with the following explanation using Phi 4 14B: {mdframed}[backgroundcolor=light-gray, linecolor=black, linewidth=0.5pt] “The provided test method testTransform exhibits the test smell known as Assertion Roulette. This occurs because the method contains multiple assertions:
  • assertEquals("123", calendar.getId());
  • assertEquals("Calendar", calendar.getName());
Each of these assertions does not include an explanatory message or parameter that clarifies what each assertion is specifically verifying. Without messages, it becomes difficult to diagnose which particular condition failed when a test fails, as all failures will simply indicate Assertion Failed without context. … ” Llama 3.2 3B, Gemma 2 9B and DeepSeek-R1 14B also detect it.
Llama 3.2 3B achieved the best performance in detecting all test smells, consistently identifying every instance. While Gemma 2 9B and Phi 4 14B also performed well, they missed a few instances. However, the explanations provided by Gemma 2 9B and Phi 4 14B were more detailed and accurate compared to those generated by Llama 3.2 3B.
These outcomes are highly encouraging and indicate that even smaller, open models can reliably detect test smells when provided with succinct, precise definitions. Unlike specialized static-analysis tools [4] that must support every possible testing framework, SLMs offer a semantic approach. We only need to provide the test smell definition in natural language. It can recognize assertions written in Mockito or JUnit without requiring an exhaustive list of method signatures to be examined.
@@ 4.2 Refactoring Test Smells
The optimal configuration for Llama 3.2 3B and Gemma 2 9B involved using four agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively. DeepSeek-R1 14B achieved its best performance with a single agent, correctly refactoring 39.3% instances. In contrast, Phi 4 14B performed best with two agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup.
Using single, two, or four agents with Llama 3.2 3B, Gemma 2 9B, or Phi 4 14B, correctly remove the Exception Handling test smell. They produce the code shown in Listing 1, where the try-catch block is removed, and the fail statement is replaced with an assertThrows statement.
Listing 1: Correct refactoring for removing Exception Handling using one, two and four agents with Gemma 2 9B.
@@
In some cases, using one agent with Gemma 2 9B to refactor results in incorrect code that alters the program’s behavior by adding an extra assertion when removing the Conditional Test Logic test smell. In contrast, using two or four agents produces a correct refactoring, effectively separating the logic into two methods.
@@
We can also correctly apply a refactoring to remove Duplicate Assert test smells using two or four agents with Phi 4 14B. However, using a single agent with Phi 4 14B applies a refactoring strategy that differs from the one proposed for this smell. Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
@@
DeepSeek-R1 14B provides significantly more details about its reasoning process compared to the other models. In some cases, it initially proposes a valid refactored code or a correct refactoring strategy during its thought process but later revises its approach, ultimately leading to an incorrect solution.
@@
For the Assertion Roulette test smell, Phi 4 14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single agent. In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies. For Duplicate Assert, the highest success rate (70%) was obtained by Phi 4 14B using two agents. Phi 4 14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single agent. The best results for Magic Number came from DeepSeek-R1 14B with two agents, achieving 90% accuracy. Overall, two- or four-agent setups led to better outcomes in three of the five test smells. However, for Assertion Roulette and Exception Handling, a single agent was sufficient. Notably, there are 23 unique instances successfully refactored by at least one model – Llama 3.2 3B (6), Gemma 2 9B (14), or DeepSeek-R1 14B (6) – that Phi 4 14B with two agents failed to refactor. By combining strengths across models, we achieve correct refactorings for 70.6% instances.
The models occasionally produced incorrect refactorings. For Assertion Roulette, common issues included misplaced messages (e.g., using JUnit 4 syntax in JUnit 5) or invalid constructs (e.g., withMessage instead of assertWithMessage). In Magic Number, the model often extracted some constants but missed others, or failed to initialize newly introduced variables. For Conditional Test Logic and Duplicate Assert, errors frequently resulted from subtle logic changes or disagreements among agents regarding whether the smell had been fully resolved. In Exception Handling, some refactorings inadvertently removed critical code along with the try block. Additionally, disagreements between Agent 3 (refactoring) and Agent 4 (evaluation) occasionally led to valid transformations being rejected—an issue observed across multiple smell types.
@@
To improve performance under the default configuration, we conducted an experiment to assess the impact of temperature [19] settings on Phi 4 14B’s ability to detect and refactor test smells. A temperature of 0.9 produced the best results for both tasks. Additionally, we allowed Phi 4 14B with four agents to make up to five attempts, rather than limiting it to just one. With pass@5 [20], it successfully detected 96% and refactored 75.3% of the test smell instances.

### v2

# Agentic LMs: Hunting Down Test Smells

@@ Abstract
Test smells reduce test suite reliability and complicate maintenance. While many methods detect test smells, few support automated removal, and most rely on static analysis or machine learning. This study evaluates models with relatively small parameter counts—Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B—for their ability to detect and refactor test smells using agent-based workflows. We assess workflows with one, two, and four agents over 150 instances of 5 common smells from real-world Java projects. Our approach generalizes to Python, Golang, and JavaScript. All models detected nearly all instances, with Phi-4-14B achieving the best refactoring accuracy (pass@5 of 75.3%). Phi-4-14B with four-agents performed within 5% of proprietary LLMs (single-agent). Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, one agent sufficed. We submitted pull requests with Phi-4-14B-generated code to open-source projects and six were merged.
@@ 1 Introduction
Test smells are design flaws in test code that compromise reliability and hinder maintenance [1]. They can reduce the effectiveness of test suites. These issues are prevalent in both open-source and industrial settings [2]. Common test smells include undocumented assertions (Assertion Roulette), complex conditional logic (Conditional Test Logic), duplicated assertions (Duplicate Assert), and hard-coded values (Magic Number[2].
Existing tools for detecting test smells typically rely on static analysis [3] or machine learning [4, 3], but these approaches are often hard to adapt to new smells or other languages [3]. Recent advances in foundation models have opened new possibilities in software engineering, yet their use in test smell refactoring – especially through collaborative agentic workflows – remains underexplored.
In this work, we evaluate open models (Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B) for automatically detecting and refactoring test smells through agentic workflows. Our method supports one, two, or four agents and is easily extensible: new smells and refactorings can be defined in natural language. We evaluate 150 instances of 5 frequent test smells from real-world Java projects (Section 3).
Results show that all models detected nearly all test smell instances (pass@5 of 96% with four-agents), with Phi-4-14B achieving the highest refactoring accuracy (pass@5 of 75.3%). This performance is within 5% of proprietary LLMs using a single-agent. Multi-agent setups outperformed single-agent ones in three of five smell types, though for Assertion Roulette, a single-agent configuration proved more effective.
Preliminary results demonstrate promising generalization across programming languages – specifically Python, Golang, and JavaScript – using the same setup. Beyond automating the detection and refactoring of test smells, our approach also enables developers to interact with models to better understand the rationale behind suggested changes. All artifacts are publicly available online. [^1] https://zenodo.org/records/16740114
@@ 2 Test Smells
Test smells are recurring patterns that reduce test clarity and maintainability. Assertion Roulette [1] occurs when a test method contains more than one assertion statement without an explanation or message (parameter in the assertion method). To mitigate this smell, developers should add a message to each assertion. Conditional Test Logic [5] arises when a test method contains one or more control statements (i.e., conditional expression, and loop statements). This can be addressed by having the developer create a dedicated test method for each condition. Duplicate Assert [2] appears when a test method contains more than one assertion statement with the same parameters. To address this, developers should split assertions that test different scenarios or states into separate tests for clarity. Exception Handling occurs when a test method contains either a throw statement or at least a catch clause. To avoid this smell, use the testing framework’s features (e.g., assertThrows) instead of manually catching or throwing exceptions. Magic Number occurs when a test method contains an assertion with a numeric literal as an argument. Refactoring involves extracting and initializing all magic numbers into constants or local variables with descriptive names.
@@ 3 Methodology
The primary goal of this study is to evaluate the effectiveness of an agentic approach using models to detect and remove test smells. We analyzed test methods from 11 real-world open-source GitHub projects using JUnit-5 previously studied by Soares et al. [6], including janusgraph, quarkus, testcontainers-java, opengrok, jenkins, lettuce, Mindustry, data-transfer-project, Activiti, flowable-engine, and skywalking. We limited our analysis to tests with up to 30 LOC, as 89% of test cases in the studied projects fall within this range. We focused on 5 common test smell types [2], selecting 30 real-world examples for each. We evaluate smaller models up to 14B parameters: Llama-3.2-3B, Gemma-2-9B, DeepSeek-R1-14B, and Phi-4-14B using their default configurations. All models were accessed via the Ollama platform and executed locally on a MacBook Pro M3 with 18GB of RAM (July 2025).
Agent communication is managed using the LangChain API, and we apply prompting techniques such as Role (Persona) and Chain-of-Thought [7] to enhance reasoning and contextual understanding. The four-agent configuration (Figure 1) distributes responsibilities as follows: Agent1 detects potential smells; Agent2 confirms the detection; Agent3 performs the refactoring; and Agent4 verifies the correctness of the refactored code. If there is disagreement, agents engage in an Evaluator-Optimizer loop [^2] https://www.anthropic.com/research/building-effective-agents, repeated up to three times.
@@
Agent1 utilizes the following prompt to detect test smells. The prompts use variables to represent relevant information required for each analysis. For instance, test_smell_definition_and_refactoring comprises two sentences: one that defines the smell and another that suggests an appropriate refactoring (see Section 2 for example sentences). You are a coding assistant with many years of experience that detects test smells.
test_smell_definition_and_refactoring
@@
You are a coding expert reviewing the detection of a test smell. Consider the following test smell.
test_smell_definition_and_refactoring
@@
It gave the following answer.
agent_answer
@@
To remove a test smell, Agent3 utilizes the prompt outlined next. During the second or third iteration, it incorporates feedback from Agent4. You are a coding assistant specializing in test code analysis and refactoring, with many years of experience.
test_smell_definition_and_refactoring
@@
To evaluate the refactored_code proposed by Agent3, Agent4 utilizes the following prompt.
You are a code reviewer specializing in JUnit 5 test smells.
test_smell_definition_and_refactoring
@@
The two-agent setup follows the same workflow but omits Agents2,4. Agent1 handles detection, and Agent3 performs the refactoring, without validation or feedback loops. In the single-agent setup, Agent1 analyzes the test code, identifies potential smells, and applies the corresponding refactoring.
@@ 4 Evaluation
Table I summarizes the detailed performance (pass@1) of the single-agent and multi-agent approaches for detecting and refactoring test smells.
Table I: Detection and refactoring performance (pass@1) of test smells by one, two, and four agents using Llama-3.2-3B (Lla), Gemma-2-9B (Gem), DeepSeek-R1-14B (DS) and Phi-4-14B (Phi).
@@
Cond. Test Logic 30 3.3% 6.7% 6.7% 0.0% 96.7% 93.3% 100.0% 100.0% 13.3% 0.0% 13.3% 0.0% 96.7% 93.3% 100.0% 100.0% 10.0% 3.3% 16.7% 3.3%
@@
Exception Handling 30 43.3% 13.3% 40.0% 36.7% 90.0% 50.0% 93.3% 100.0% 43.3% 23.3% 43.3% 30.0% 86.7% 50.0% 90.0% 100.0% 40.0% 23.3% 40.0% 36.7%
Magic Number 30 73.3% 83.3% 30.0% 33.3% 93.3% 100.0% 100.0% 100.0% 80.0% 90.0% 43.3% 40.0% 90.0% 100.0% 100.0% 100.0% 70.0% 90.0% 50.0% 33.3%
@@ 4.1 Detecting Test Smells
Gemma-2-9B and Phi-4-14B exhibit good detection performance across most test smells, correctly identifying 98% and 96% instances, respectively, when using two agents. Their performance remains consistent with one or four agents. For every test smell, at least one model achieves a perfect detection score. DeepSeek-R1-14B demonstrates the weakest performance, detecting only 78% instances. In contrast, Llama-3.2-3B consistently detects all test smells across all agent configurations.
We review Agents1,2 outputs to assess the validity of their test smell explanations. Next, Agent1 uses Phi-4-14B to identify the Exception Handling test smell. The other models also detect it. “The test method testStackBlowOut includes a try-catch block to handle exceptions manually, which is indicative of the Exception Handling test smell. The code attempts to catch an IllegalArgumentException, rather than using testing framework features like assertThrows to verify that the exception is thrown as expected.” Llama-3.2-3B achieved the highest detection rate, identifying all test smell instances. Gemma-2-9B and Phi-4-14B missed a few but provided more detailed and accurate explanations. These results demonstrate that smaller models can reliably detect test smells when given concise natural language definitions. Unlike static-analysis tools, models take a semantic approach, recognizing frameworks like Mockito or JUnit without needing exhaustive method lists.
@@ 4.2 Refactoring Test Smells
The optimal configuration for Llama-3.2-3B and Gemma-2-9B involved using four-agents, successfully refactoring 22% and 40.7% out of 150 instances in one attempt, respectively. DeepSeek-R1-14B achieved its best performance with a single-agent, correctly refactoring 39.3% instances. In contrast, Phi-4-14B performed best with two-agents, successfully refactoring 55.3% instances, demonstrating its effectiveness in a multi-agent setup. Processing 150 smells using four-agents took ~1h (Llama-3.2-3B, Gemma-2-9B), ~5h (Phi-4-14B), and ~8h (DeepSeek-R1-14B).
Using single, two, or four-agents with Llama-3.2-3B, Gemma-2-9B, or Phi-4-14B, correctly remove the Exception Handling test smell. For instance, they produce the code shown in Listing 1, where the try-catch block is removed, and the fail statement is replaced with an assertThrows statement.
Listing 1: Correct refactoring for removing Exception Handling using single and multi-agents with Gemma-2-9B.
@@
In some cases, using single-agent with Gemma-2-9B to refactor results in incorrect code that alters the program’s behavior by adding an extra assertion when removing the Conditional Test Logic test smell. In contrast, using multi-agents produces a correct refactoring, effectively separating the logic into two methods.
@@
We can also correctly apply a refactoring to remove Duplicate Assert test smells using multi-agents with Phi-4-14B. However, using a single-agent with Phi-4-14B applies a refactoring strategy that differs from the one proposed for this smell. Specifically, it incorporates a unique message for each assertion, effectively eliminating both this test smell and the Assertion Roulette test smell.
@@
DeepSeek-R1-14B provides significantly more details about its reasoning process compared to the other models. In some cases, it initially proposes a valid refactored code or a correct refactoring strategy during its thought process but later revises its approach, ultimately leading to an incorrect solution.
@@
For the Assertion Roulette test smell, Phi-4-14B achieved the best result, correctly detecting and refactoring 93.3% of instances with a single-agent. In contrast, all models struggled with Conditional Test Logic, where we evaluated only one of several possible refactoring strategies. For Duplicate Assert, the highest success rate (70%) was obtained by Phi-4-14B using two-agents. Phi-4-14B also performed best on Exception Handling, correctly refactoring 43.3% of cases with a single-agent. The best results for Magic Number came from DeepSeek-R1-14B with two agents, achieving 90% accuracy. Overall, two or four-agent setups led to better outcomes in three of the five test smells. However, for Assertion Roulette and Exception Handling, a single-agent was sufficient. Notably, there are 23 unique instances successfully refactored by at least one model – Llama-3.2-3B (6), Gemma-2-9B (14), or DeepSeek-R1-14B (6) – that Phi-4-14B with two-agents failed to refactor. By combining strengths across models, we achieve correct refactorings for 70.6% instances (pass@1).
@@
To improve performance under the default configuration, we conducted an experiment to assess the impact of temperature [8] settings on Phi-4-14B’s ability to detect and refactor test smells. A temperature of 0.9 yielded the best results for both tasks, slightly improving performance by 3.3%.
4.3 Failure Types
The ablation study revealed distinct failure patterns across agent configurations. In the single-agent setup, issues frequently included incomplete fixes, superficial code changes that failed to resolve the smell, and the introduction of invalid or non-existent constructs—often stemming from incorrect syntax or reliance on frameworks other than JUnit-5. In Magic Number, the model sometimes extracted only a subset of constants or failed to initialize new variables. Refactorings for Conditional Test Logic and Duplicate Assert occasionally introduced subtle logic changes. In Exception Handling, some transformations mistakenly removed critical code alongside the try block.
Introducing additional agents reduced some errors but brought new challenges, including increased code divergence, missed detections that the single agent previously succeeded on, and frequent disagreements—particularly with Agent4 rejecting correct refactorings proposed by Agent3. Moreover, if the multi-agent setup had used JUnit-5 instead of JUnit-4 and correctly positioned the message parameter, its performance in Assertion Roulette would have matched that of the single-agent configuration.